# Early-Stage Anomaly Detection: A Study of Model Performance on Complete vs. Partial Flows

### v1

@@ IV-A Packet Count-based Evaluation
Figure 1: Comparative analysis of binary and multi-class classification performance across different packet count scenarios.
IV-A1 CF and PF Dataset Distribution
Table IV presents the dataset distributions for packet count-based evaluation. The table starts with the ”Complete Flows” (CF) category, which encompasses the ground truth dataset (see Section III-D) and then breaks down into partial flows labeled as PC=N, where N ranges from 2 to 20, representing datasets characterized by specific packet counts. For each category, Table IV lists the total number of flows, classified into benign and anomalous types, with anomalies further subdivided into distinct types of DoS attacks. The table also includes metrics on the minimum, mean, and maximum durations (in milliseconds) for these categories.
@@ IV-A1 CF and PF Dataset Distribution
IV-A2 CF vs. PF Performance Comparison
Figure 1 illustrates the precision, F1-score, and recall for both binary (Figure 1a) and multi-class (Figure 1b) classification under three scenarios:
@@
From Figure 1a, it is observed that both CF train/CF test and N train/N test scenarios achieve sufficient performance starting at PC=2, though performance is slightly lower at minimal PC values in the N train/N test scenario. However, the CF train/N test scenario highlights the negative impacts of insufficient information due to lower packet counts, indicating that at least 8 packets are necessary in the N test to achieve usable performance using CF train model, with some minor fluctuations at higher packet counts.
The multi-class classification in Figure 1b similarly shows that low packet counts negatively affect the N train/N test scenario, requiring at least 6 packets to achieve acceptable performance. Additionally, even the highest packet count seems inadequate for stable performance across precision, F1-score, and recall in the CF train/N test scenario, underscoring the challenges of anomaly detection with limited data when only completed flows have been used for training.
IV-B Flow Duration-based Evaluation
@@
Figure 2: Comparative analysis of binary and multi-class classification performance across different flow duration scenarios.
IV-B1 CF and PF Dataset Distribution
Similar to Table IV, Table V outlines the dataset distributions for flow duration-based evaluation. This methodical categorization delineates a transition from ”Complete Flows” (CF) to progressively more granular ”Partial Flows” (PF), identified as FD=N, where N encompasses a range of specified durations ${5,10,50,100,150,300,500,1000,5000,10000,15000,20000}$ milliseconds. Each category within Table V specifies the total number of flows, distinguishing between benign and anomalous flows, with the latter further divided into distinct DoS attack types. Contrasting with Table IV, which detailed flow durations, this table provides insights on the minimum, mean, and maximum packet counts for these categories.
Upon examining Table V, a similar trend emerges as was seen with packet count-based datasets; however, when flow duration thresholds define partial flow generation, there’s a dramatic drop in flow counts for each dataset. This decrease persists even though our methodology accommodates flows with durations within $\pm$20of the threshold values, as detailed in Section III-E.
@@ IV-B1 CF and PF Dataset Distribution
IV-B2 CF vs. PF Performance Comparison
Figure 2 depicts the performance metrics of precision, F1-score, and recall for binary and multi-class classifications across a range of FD thresholds. For the sake of this analysis, only FD=5, 10, 150, 500, 1000, 15000, and 20000 ms are included, as other durations did not have enough data to yield statistically significant results. Analogous to the packet count-based evaluation, prior to training, we exclusively selected flows from the CF dataset that correspond with the flow hash identifiers present within the respective partial flow dataset.
Figure 2a confirms the trend observed in packet count evaluations. When the training and testing datasets match (CF train/CF test and N train/N test), binary classification performance is strong and stable. However, the model’s ability to accurately classify benign and anomalous flows is compromised in the CF train/N test scenario, reflecting the challenge of applying a model trained on complete flow data to partial flows characterized by various durations.
In Figure 2b, this issue is further exacerbated in the multi-class classification scenario. Here, even the N train/N test performance demonstrates a notable decline, particularly at FD=1000ms, suggesting that as with certain flow durations set as threshold, the model’s ability to distinguish between multiple classes of traffic—beyond benign and anomalous—becomes is unreliable. This decline in performance may be attributed to the reduced information content in flows with various duration, which does not adequately capture the complexities of multi-class traffic patterns.
@@
V Discussion
This study expands our packet count- and flow duration-based evaluations with additional digital artifacts [github-repo]. These materials provide a more comprehensive set of metrics, including accuracy, balanced accuracy, and confusion matrices. Moreover, they offer visualizations that extend beyond the scope of those included in the main body of this paper.
V-A Implications
Our findings confirm that models trained and tested on the same data distribution (CF or N) exhibit commendable performance. However, a decrement in packet count introduces a performance dip, suggesting that a model with less data may struggle to generalize well. This effect is particularly pronounced in cases where models trained on complete data (CF) are applied to partial data (PC=N), highlighting potential overfitting to the complete data’s characteristics.
The analysis indicates that while higher packet counts are associated with extended flow durations, attacks with fewer packets could present challenges for precise anomaly detection due to insufficient detail in flow statistics. For effective binary classification, a threshold of at least 7 packets in the test data is essential to ensure performance reliability when classifying using CF dataset (CF train/N test), while for multi-class classification, at least 6 packets are necessary when trained and tested on the same dataset (N train/N test).
Fluctuations in performance across various packet counts emphasize the significance of selecting an optimal packet count threshold for maintaining classification accuracy. When training and testing datasets are aligned, performance remains stable. Yet, the application of models developed on complete flows to partial flows poses a significant challenge, underscoring the importance of data quality and volume in training robust anomaly detection models.
Additionally, the non-linear trends observed across different flow duration thresholds highlight the complexity inherent in network attack profiling, which cannot be solely based on duration metrics. The adaptive and dynamic nature of such attacks necessitates a more nuanced detection approach that can respond to varied attack patterns and evasion techniques.
Interestingly, while a specific packet count threshold can be identified that yields stable classification performance, no such threshold emerged for flow duration. This underscores the challenge in defining a one-size-fits-all threshold for flow duration-based classification.
Overall, the research presented here sheds light on the complex interplay between data distribution and classification efficacy, leading to essential considerations for deploying network security measures. Future models must be designed with the flexibility to adjust to varying levels of traffic information, ensuring that anomaly detection remains both robust and reliable.
V-B Limitations
While this research provides valuable insights into anomaly detection using Random Forest models, it is not without limitations that may affect the generalizability and application of the findings. First, the study does not extend to the examination of flows with variable lengths. Our analysis primarily focuses on flows characterized by a fixed number of packets, assessing them from both packet count and flow duration perspectives. This approach limits our ability to fully understand the dynamics of more complex, variable-length network flows that often occur in real-world environments. The fixed-length assessment may overlook nuances that appear only in longer or dynamically changing sequences, potentially affecting the model’s ability to generalize across different network conditions.
Secondly, the current study focuses exclusively on the first $N$ packets per flow. This methodology may miss crucial information contained in subsequent packets of a flow, especially in longer or continuous data streams. It might be beneficial to assess the flows through a sliding window approach or a sequence of sub-flows, which could capture changes and anomalies that develop beyond the initial packets.
Additionally, analyzing partial flows with packet counts starting from 2 may not adequately address the complexities of modern protocols such as TLS and QUIC, where connection or session setup alone may require more packets than the lower range of our study. Focusing on such an early packet range might overlook crucial information for classification that appears only during or after the completion of these setups.
Packet count thresholds also face challenges with hardware offloading technologies, which, after a certain number of packets, can shift the maintenance of flows to fast path processing. Such hardware-level interventions can obscure or completely hide parts of the flow from observation, complicating detection strategies that rely on early-stage categorization.
Furthemore, our method of generating partial flows based on flow duration with different thresholds highlights a challenge. While allowing a degree of flexibility by retaining flows whose durations fall within a ±$\pm 20\%$ range of a specified target duration, a dramatic drop in flow counts for each dataset indicates that our approach may still be too restrictive. A more refined methodology is needed to better assess and accommodate the variable nature of flow durations, ensuring that our anomaly detection framework can effectively handle the diverse and dynamic characteristics of network traffic.
Lastly, this study examines the performance of the RF algorithm. To assess generalizability, we also evaluated Decision Trees (DT), whose codes are accessible via [github-repo]. While DT approached the performance of RF, generally, RF outperformed DT in our evaluations. However, an extended analysis involving a broader range of various MLAs is needed to draw more targeted conclusions about the effectiveness of different models.
V-C Related Work
@@
While these studies contribute significantly to the field, they often do not comprehensively examine crucial aspects such as packet count and flow duration, both essential for ensuring robust and reliable anomaly detection in real-world scenarios. Our research contrasts the use of packet count and flow duration thresholds for early-stage anomaly detection, establishing a foundational understanding of the trade-offs between various threshold strategies for network traffic analysis. This nuanced approach allows our study to provide practical insights that can significantly influence real-time network security measures.
V-D Comparison with Relevant Research
@@
VI Conclusion
This study examined the efficacy of the Random Forest model in the context of anomaly detection, applying it to both complete and partial network flow data derived from a refined version of the CICIDS2017 dataset. Our findings underscore the necessity of using at least 7 packets to achieve reliable anomaly detection performance, contrasting with recent studies that suggest fewer packets may be sufficient.
Our research contributes to the development of machine learning models that accurately reflect the challenges of analyzing partial flow data in real-time network environments. By aligning model development and validation with the operational dynamics inherent in real-time anomaly detection, this study helps bridge critical gaps in current research methodologies.
Future work will focus on refining data preprocessing techniques to handle variable-length data more effectively and exploring the integration of more complex machine learning models that can automatically adjust to the dynamic nature of network traffic, potentially improving the accuracy and responsiveness of anomaly detection systems.
@@ Acknowledgement
This work was supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. Supported by the ÚNKP-23-5-BME-461 New National Excellence Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund. The work presented in this paper was supported by project no. TKP2021-NVA-02. Project no. TKP2021-NVA-02 has been implemented with the support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme.

### v3

@@ 4.1 Packet Count-based Evaluation
Figure 1: Comparative analysis of binary and multi-class classification performance across different packet counts.
4.1.1 CF and PF Dataset Distribution
Table IV presents the dataset distributions for packet count-based evaluation. The table starts with the Complete Flows (CF) category, which encompasses the ground truth dataset (see Section 3.4) and then breaks down into partial flows labeled as PC=N, where N ranges from 2 to 20, representing datasets characterized by specific packet counts. For each category, Table IV lists the total number of flows, classified into benign and anomalous types, with anomalies further subdivided into distinct types of DoS attacks. The table also includes metrics on the minimum, mean, and maximum durations (in milliseconds) for these categories.
@@ 4.1.1 CF and PF Dataset Distribution
4.1.2 CF vs. PF Performance Comparison
Fig. 1 illustrates the precision, F1-score, and recall for both binary (Fig. 1a) and multi-class (Fig. 1b) classification under three scenarios:
@@
From Fig. 1a, it is observed that both CF train/CF test and N train/N test scenarios achieve sufficient performance starting at PC=2, though performance is slightly lower at minimal PC values in the N train/N test scenario. However, the CF train/N test scenario highlights the negative impacts of insufficient information due to lower packet counts, indicating that at least 8 packets are necessary in the N test to achieve usable performance using CF train model, with some minor fluctuations at higher packet counts.
The multi-class classification in Fig. 1b similarly shows that low packet counts negatively affect the N train/N test scenario, requiring at least 6 packets to achieve acceptable performance. Additionally, even the highest packet count seems inadequate for stable performance across precision, F1-score, and recall in the CF train/N test scenario, underscoring the challenges of anomaly detection with limited data when only completed flows have been used for training.
The results show that conclusions drawn from evaluations using only complete flows may not generalize to operational environments. Notably, while previous works have reported high detection rates using this dataset, our findings suggest that such performance metrics should be interpreted with caution when considering real-time detection scenarios.
4.2 Flow Duration-based Evaluation
Next, we analyze scenarios where partial flows are defined by reaching a specific flow duration ($N_{fd}$) threshold.
@@ 4.2 Flow Duration-based Evaluation
Figure 2: Comparative analysis of binary and multi-class classification performance across different flow durations
4.2.1 CF and PF Dataset Distribution
Similar to Table IV, Table V outlines the dataset distributions for flow duration-based evaluation. This methodical categorization delineates a transition from Complete Flows (CF) to progressively more granular Partial Flows (PF), identified as FD=N, where N encompasses a range of specified durations $5,10,50,100,150,300,500,1000,5000,10000,15000,$ $20000$ milliseconds. Each category within Table V specifies the total number of flows, distinguishing between benign and anomalous flows, with the latter further divided into distinct DoS attack types. Contrasting with Table IV, which detailed flow durations, this table provides insights on the minimum, mean, and maximum packet counts for these categories.
Upon examining Table V, a similar trend emerges as was seen with packet count-based datasets; however, when flow duration thresholds define partial flow generation, there is a dramatic drop in flow counts for each dataset. This decrease persists even though our methodology accommodates flows with durations within $\pm$20of the threshold values, as detailed in Section 3.5.
@@ 4.2.1 CF and PF Dataset Distribution
4.2.2 CF vs. PF Performance Comparison
Fig. 2b depicts the performance metrics of precision, F1-score, and recall for binary and multi-class classifications across a range of FD thresholds. For the sake of this analysis, only FD=5, 10, 150, 500, 1000, 15000, and 20000 ms are included, as other durations did not have enough data to yield statistically significant results. Analogous to the packet count-based evaluation, prior to training, we exclusively selected flows from the CF dataset that correspond with the flow hash identifiers present within the respective partial flow dataset.
Fig. 2a confirms the trend observed in packet count evaluations. When the training and testing datasets match (CF train/CF test and N train/N test), binary classification performance is strong and stable. However, the model’s ability to accurately classify benign and anomalous flows is compromised in the CF train/N test scenario, reflecting the challenge of applying a model trained on complete flow data to partial flows characterized by various durations.
In Fig. 2b, this issue is further exacerbated in the multi-class classification scenario. Here, even the N train/N test performance demonstrates a notable decline, particularly at FD=1000ms, suggesting that as with certain flow durations set as threshold, the model’s ability to distinguish between multiple classes of traffic—beyond benign and anomalous—becomes is unreliable. This decline in performance may be attributed to the reduced information content in flows with various duration, which does not adequately capture the complexities of multi-class traffic patterns.
@@
The irregular trends observed in flow counts and performance metrics highlight a critical aspect often overlooked in previous evaluations: the complex relationship between flow duration and detection capability. This relationship has significant implications for how we interpret past studies that assumed complete flow information would be available.
5 Discussion
Our empirical results provide several critical insights into the performance of ML-based network anomaly detection when faced with the reality of partial flow information. This discussion synthesizes these findings, considers their implications, and outlines limitations and future research avenues.
5.1 Implications of Findings
The systematic comparison across complete and partial flow scenarios reveals significant implications. Our primary finding is the quantifiable performance degradation (up to 30% drop in precision/recall in some scenarios) when models trained on CF are tested on PF, particularly those defined by low packet counts or short durations (Fig. 1, Fig. 2b). This directly challenges the implicit assumption in many studies that high performance achieved using CF translates directly to real-time operational effectiveness. Practitioners should be cautious when interpreting lab-based results derived solely from CF.
Next, the study highlights differing sensitivities to threshold types. While consistent performance was achievable in packet count scenarios (PC=N Train/PC=N Test) above a certain threshold (around 7 packets for the CF Train/PF Test binary case), performance based on flow duration thresholds was more volatile (Table V, Fig. 2b). This suggests packet count might offer a more stable trigger for early analysis than fixed time windows for this dataset and model, although both require careful consideration.
The observation that approximately 7-8 packets were needed in the test set for a CF-trained RF model to achieve reasonable binary classification performance (Section 4.1.2) provides a concrete, albeit dataset-specific, benchmark for the minimum information required under these mismatched conditions. This finding has direct implications for designing early detection systems.
The results also underscore the inherent trade-off between detection speed (requiring early decisions on partial data) and accuracy/reliability (which often improves with more complete data). The poor performance in the CF Train/PF Test multi-class scenarios further emphasizes that complex discriminations require richer information, which may not be available early in a flow’s lifecycle.
Lastly, the common practice of evaluating NIDS primarily on complete flow datasets may provide an overly optimistic view of real-world capability. Our work advocates for evaluation methodologies that explicitly incorporate partial flow scenarios, particularly the CF Train/PF Test mismatch, to better assess practical deployability.
5.2 Limitations and Future Directions
While this study provides valuable quantitative insights, certain limitations define avenues for future research.
Evasion Potential: The finding that around 7 packets were needed for reliable detection by the CF-trained model (Section 4.1.2) highlights a potential vulnerability. Sophisticated attackers could potentially evade such early detection by mimicking benign behavior for the initial packets before launching malicious activity. Addressing this requires further research into adaptive detection mechanisms, analysis of feature evolution within flows over time, robust feature engineering less susceptible to mimicry, and potentially combining early detection with later-stage analysis as part of a defense-in-depth strategy. This vulnerability underscores that reliance solely on very early detection can be risky.
Generalizability—Dataset: Our analysis is based on a single dataset (CICIDS-2017, Wednesday traffic). While chosen for its prevalence and availability of packet data, its age (2017 traffic) and known issues [engelen2021, lanvin2023, Flood2024, pekar2024evaluating] mean the findings require validation on newer, more diverse datasets reflecting contemporary traffic patterns, attack vectors, and network conditions (e.g., varying baseline traffic, different protocols). Real-world network traces would provide the ultimate test.
Generalizability—Algorithm: We focused on Random Forest as a representative and widely used algorithm. While this provided a clear baseline, the extent to which these specific quantitative results (e.g., the 7-packet threshold) generalize to other ML algorithms (e.g., SVM, Deep Learning models like LSTMs or CNNs, or unsupervised methods like Isolation Forest or Autoencoders) is an open question. Future work should compare different algorithmic approaches under the same partial flow constraints.
Study Assumptions Rationale: Examining the CF Train / PF Test scenario was deliberate. It directly models the common situation where research prototypes (often CF-trained) are considered for deployment in environments dealing with PFs. Quantifying this gap is crucial for practitioners. While training specialized models for different packet counts/durations might seem ideal, the unpredictable nature of real-world flows makes maintaining numerous specialized models challenging. Our study informs the design of single, more robust models tolerant to varying completeness. Hybrid approaches or adaptive models remain interesting future directions.
Variable-length Flows & Feature Evolution: Our fixed-threshold approach (PC=N, FD=N) simplifies analysis but does not fully capture the dynamic evolution of features within a single flow over time. Future work could explore methods that analyze sequences of partial flow updates or use techniques sensitive to feature changes as a flow progresses.
Protocol Complexities: Modern encrypted protocols (TLS 1.3, QUIC) often require more packets for handshake completion than our lower packet count thresholds (e.g., PC=2). Analyzing the impact of partial information specifically within the context of these complex handshakes and application data phases is needed.
Hardware Offloading Impact: Hardware offloading can reduce the visibility of flows after an initial packet count, posing challenges for continuous monitoring. Investigating detection techniques robust to such hardware limitations is warranted.
Flow Duration Methodology Refinement: The volatility observed in datasets generated by duration thresholds, even with flexibility, suggests that more sophisticated time-based sampling or feature extraction methods might be needed for robust duration-based early detection.
These limitations highlight opportunities for refinement and extension, building upon the foundation established here regarding the critical role of data completeness in anomaly detection performance.
6 Related Work
Our work intersects with research on early-stage network traffic analysis and studies specifically evaluating NIDS performance under varying conditions.
6.1 Early-Stage Network Traffic Analysis
@@
While valuable, these studies often focus on a single mechanism (e.g., time windows, fixed packet count) or algorithm. Our work provides a direct, systematic comparison of performance across three distinct train/test paradigms (CF/CF, PF/PF, CF/PF) using both packet count and flow duration as triggers for partial flow generation. The key contribution is the quantification of the performance impact specifically due to the train/test mismatch using these common triggers, providing a nuanced understanding relevant to practical deployment trade-offs.
6.2 Comparison with Relevant Research
@@
7 Conclusion
This study presents a systematic reappraisal of how partial flow information affects machine learning-based anomaly detection. Using a refined version of the CICIDS-2017 dataset, we demonstrate that at least 7 packets are necessary for reliable detection, contrasting with recent studies suggesting fewer packets may suffice. Our findings have important implications for how we interpret previous evaluations using this dataset, particularly regarding the gap between laboratory results and operational requirements. The methodology and insights presented here contribute to better evaluation practices for network security tools. By explicitly considering the impact of partial flows, we highlight a critical aspect often overlooked in previous studies. This work also demonstrates the importance of carefully examining assumptions about data completeness in security evaluations.
8 Reproducibility and Extended Analysis
To support reproducibility and facilitate further investigation of our findings, a digital artifact containing all necessary data and code is publicly available at [github-repo]. This includes the carefully processed datasets used for training and testing (aligned with the methodologies in Section 3), the analysis scripts used to generate the results, and performance metrics. Beyond the core metrics presented in Section 4, the artifact also provides extended analyses, and complementary visualizations, offering deeper insights into specific class performance and error types.
@@ Acknowledgement
Supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. This work was also supported by projects TKP2021-NVA-02 (financed under the TKP2021-NVA scheme) and 2024-1.2.6-EUREKA-2024-00009 (financed under the 2024-1.2.6-EUREKA scheme), implemented with support provided by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation Fund.